OpenStudy (funsize67):

Which measurements of the side of a right triangle A.202, 198,40 B.28, 26, 12 C.17,14,6 D. all of the above

OpenStudy (luigi0210):

A right triangle has the following properties \(\large a^2+b^2=c^2 \) Plug in the numbers and see if the equation is true

OpenStudy (luigi0210):

For example A \(\Large 40^2+198^2=202^2 \) \(\Large 1600+39204 = 40804 \) \(\Large 40804=40804 \) the equation turns out to be true, so A would be a right triangle

OpenStudy (funsize67):

Thanks

OpenStudy (luigi0210):

Check the remaining choices, they may be true too

OpenStudy (mrnood):

Hint: c in the equation is always the longest side from the 3 given